OpenStudy (jazzyfa30):

HCl+NaOH=H2O +NaCl Instructions For each chemical statements written below do an atom inventory for each one . Determine if it is a ablanced equation .help please

OpenStudy (twopointinfinity):

THC

OpenStudy (jazzyfa30):

??????

OpenStudy (anonymous):

Do you want to know how

OpenStudy (twopointinfinity):

\[T=Hci+Na H*O/O2 \frac{ THC }{ H2O }\]

OpenStudy (jazzyfa30):

This is the Queston is it Balanced??? Instructions Do an atom inventory for each one.

OpenStudy (twopointinfinity):

THC over H20 That is very balanced. Yes.

OpenStudy (anonymous):

Use a Elemental scale, it will help you, you can google one if you do not have one.

OpenStudy (jazzyfa30):

where did the T come from

OpenStudy (twopointinfinity):

I + N = T

OpenStudy (jazzyfa30):

???????

OpenStudy (twopointinfinity):

I'm in my 3rd year of college just believe me. If you show that to your professor he will be so impressed with you.

OpenStudy (jazzyfa30):

im not in college im in hs

OpenStudy (twopointinfinity):

Im in college though, that's the answer.

OpenStudy (jazzyfa30):

ok can you help with some more

OpenStudy (twopointinfinity):

sure. post it in another question.

OpenStudy (jazzyfa30):

KNO3=KNO2+O2

OpenStudy (twopointinfinity):

you have to post it in another question.

OpenStudy (jazzyfa30):

ok

OpenStudy (amistre64):

HCl+NaOH=H2O +NaCl ^ ^ ^ 2 Hs on each side Cl+NaO=O +NaCl ^ ^ 1 Cl on each side NaO=O +Na ^ ^ 1 Na on each side O = O ^ ^ 1 O on each side .... each side has the same number of elemetals